OpenStudy (anonymous):

whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) 2x − 6

OpenStudy (anonymous):

yes the left expression is equal to the right

jimthompson5910 (jim_thompson5910):

you can think of 2(x − 3) as (x-3) + (x-3) (x-3) + (x-3) x-3 + x-3 (x+x)+(-3-3) 2x - 6

jimthompson5910 (jim_thompson5910):

or you can distribute 2*(x-3) = 2*x - 2*3 2*(x-3) = 2x - 6 either way, the two sides are the same